The sum of the ages of A and B is 41 years now. 10 years ago, A was twice as old as B was then. Find their ages now.
Answer by ankor@dixie-net.com(22740)

The sum of the ages of A and B is 41 years now.
a + b = 41
10 years ago, A was twice as old as B was then.
a-10 = 2(b-10)
a - 10 = 2b - 20
a = 2b - 20 + 10
a = 2b - 10
Find their ages now.
In the first equation, replace a with (2b-10)
2b - 10 + b = 41
3b = 41 + 10
3b = 51
b = 51/3
b = 17 yrs
:
Find a
a = 2(17) - 10
a = 34 - 10
a = 24 yrs